OpenStudy (jollypikachu300):

WILL FAN AND MEDAL 3 - when y = -4

OpenStudy (jollypikachu300):

@OregonDuck @just_one_last_goodbye help

OpenStudy (anonymous):

can you explaqin the Qu please

OpenStudy (jollypikachu300):

thats the question 3-y when y=-4

OpenStudy (anonymous):

what the question?

OpenStudy (jollypikachu300):

evulate each expression for the given values of the variables

OpenStudy (anonymous):

it 7, just change the y with (-4) so, +3-y be 3-(-4) =3+4 =7

OpenStudy (jollypikachu300):

thx @ArifL

OpenStudy (jollypikachu300):

another one plz

OpenStudy (anonymous):

go on

OpenStudy (jollypikachu300):

b=(-8) when b=-3

OpenStudy (anonymous):

that not a question, that 2 fact

OpenStudy (jollypikachu300):

that is question 18

OpenStudy (anonymous):

what is the question? there nothing that need to be solved there

OpenStudy (jollypikachu300):

it is just like the other question we did before this

OpenStudy (anonymous):

it sill 3-y?

OpenStudy (jollypikachu300):

oh sorry

OpenStudy (jollypikachu300):

3 - y

OpenStudy (jollypikachu300):

@ArifL come on im wating for u to anwser me

OpenStudy (jollypikachu300):

@pooja195

OpenStudy (anonymous):

b-x=-8 b=-3 =(-3)-x=-8 =-x=-5 =x=5

pooja195 (pooja195):

What is the question?

OpenStudy (jollypikachu300):

scroll up to the second question

OpenStudy (jollypikachu300):

so x = 5

pooja195 (pooja195):

I cant tell....repost the question

OpenStudy (jollypikachu300):

or no b = 5

pooja195 (pooja195):

Wat? Report your question!! :P

pooja195 (pooja195):

*repost

OpenStudy (jollypikachu300):

b - (-8) when b = -3

pooja195 (pooja195):

-3+8= ? 2 negs make a pos

pooja195 (pooja195):

So you said 5 which is correct

OpenStudy (jollypikachu300):

ok so b = 5

OpenStudy (jollypikachu300):

thx

pooja195 (pooja195):

Its not b=5 the answer is just 5

OpenStudy (jollypikachu300):

ok

OpenStudy (jollypikachu300):

i got one more

OpenStudy (jollypikachu300):

-8-p-5 when p= -5

OpenStudy (anonymous):

-8-(-5)-5 =-8+5-5 =-8

OpenStudy (jollypikachu300):

thx